find the limit lim x→0 x^2 cos 1/x.

1
Expert's answer
2022-01-13T15:52:38-0500

"\lim_{x\to0}(x^2*cos({\frac 1 x}))"

"\forall a:-1≤cos(a)≤1"

"\lim_{x\to0}x^2=0\implies \lim_{x\to0}(x^2*cos({\frac 1 x}))=0*b" where "-1≤b≤1\implies \lim_{x\to0}(x^2*cos({\frac 1 x}))=0"